(Related to Checkpoint​ 9.2)  ​(Yield to​ maturity)  The market price is ​$900 for a 16​-year bond ​($1 comma 000 par​ value) that pays 8 percent annual​ interest, but makes interest payments on a semiannual basis ​(4 percent​ semiannually). What is the​ bond's yield to​ maturity? The​ bond's yield to maturity is nothing​%. ​ (Round to two decimal​ places.)
Answer #1

Answer:

Par Value = $1,000

Market price = $900

Semiannual Interest payment = 1000 * 4% = $40

Time to maturity = 16 * 2 = 32 semiannual periods

To get bond's yield to​ maturity, we will use RATE function of excel:

= RATE (nper, pmt, pv, fv, type)

= RATE (32, 40, -900, 1000, 0)

= 4.6032%

Yield to maturity = 4.6032% * 2 = 9.21%

The​ bond's yield to maturity is 9.21%